OpenStudy (anonymous):

The two diagonals of parallelogram PQRS are 32 cm and 60 cm long. If the parallelogram is a rhombus, what is the length of side PQ?

OpenStudy (anonymous):

@BTaylor @Aqua666 anyone help! please!

OpenStudy (anonymous):

Use the formula: \[A = 1/2d _{1}d _{2}\]

OpenStudy (anonymous):

A = 1/2(d1)(d2) A = 1/2(32(60) A = 32(30) A = 960

OpenStudy (anonymous):

so new formula will be A=1/2 32*60?

OpenStudy (anonymous):

Yes.

OpenStudy (btaylor):

The diagonals of a rhombus are perpendicular, so think of a triangle that looks like this:  The diagonals would be bisected into sides of 16 and 30 cm. So, with the pythagorean theorem, 16^2 + 30^2 = PQ^2 so PQ = 34.

OpenStudy (anonymous):

oh ok! thanx so much! you didnt have to answer it for me! I like people giving me the formulas (helping me out) and I have to figure out the answer, that way I am also putting a little bit of effort! YOU GUYS ARE AMAZING! THNX SO MUCH!

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

WAIT, 960 or 34?

OpenStudy (anonymous):

actually @BTaylor was with the right answer ;)